# A Fast Algorithms for Fourier extension based on boundary interval data

### v1

@@ Abstract
This paper present a new algorithm for the computation of Fourier extension based on boundary data, which can obtain a super-algebraic convergent Fourier approximation for non-periodic functions. The algorithm calculates the extension part through boundary data and connects it with the original function to form a periodic smooth function. By testing the key parameters involved, their impact on the algorithm is clarified and the optimization setting scheme of the parameters is proposed. Compared with FFT, the algorithm only needs to increase the computational complexity by a fixed small amount.
@@
Fourier Series, Fourier Extensions, Fourier Continuation
@@ 1 Introduction
Fourier series approximation for smooth periodic functions is popular because it can offer spectral convergence and be computed numerically via the FFT. However, the situation is changed completely once periodicity is lost. A jump discontinuity is imposed in the approximation at the domain boundaries due to the well known Gibbs phenomenon. Many ways have been developed to ameliorate the ill effects of the Gibbs phenomenon, including the methods based on Gegenbauer polynomials [9, 10, 16], the methods utilizing Padé approximations [7, 8] and the Fourier extension (or continuation) methods [4, 11, 12, 14, 13, 15, 6, 17].
@@
To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system need to be solved. In [4], Boyd has introduced the truncated singular value decomposition (TSVD) to stabilize the calculation process, the accuracy almost up to machine precision. A convergence analysis by Adcock et. al states that for analytic functions the method is at least superalgebraically convergent [2]. Utilizing randomized algorithms, Lyon in [12] provided a fast algorithm for the computation with $T=2$. And in [14], Matthysen et. al presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation. This idea was further developed in the literature [15] to form a fast algorithm for function approximation on 2D domains.
It should be noted that the goal of the extension process is to effectively connect the data at both ends of the function. We construct the following two functions:
$f_{1}(t)=\left\{\begin{aligned} &-1,&x<-\frac{1}{2}\\ &\tanh\left(\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2}\\ &1,&t>\frac{1}{2}\end{aligned}\right.$ (3)
$f_{2}(t)=\left\{\begin{aligned} &-1,&t<-\frac{1}{2}\\ &\tanh\left(100*\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2}\\ &1,&t>\frac{1}{2}\end{aligned}\right.$
Since the data of the two functions near the endpoints are the same, the same connection data can be used on the interval $[-T,1)$ and $(1,T]$. An illustration is shown in Fig. 1. This raises an issue: does the data inside the interval really help the calculation of the extension? In fact, A very effective method, known as FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations [5]. Recently, this method has also been extended to the approximation of two-dimensional functions[6]. The numerical implementation of the FC-gram algorithm requires a high-precision environment and it is not superalgebraic convergent.
@@
This paper proposes a new fast algorithm for the calculation of Fourier extension, which has the following traits:
  • Only a few discretization points in the boundary interval are used to calculate the extension part, so the method can be computed with FFT speed. Unlike the FC-Gram algorithm, the method in this paper does not need to introduce polynomials.
  • Since the extension part only uses boundary data for calculation, the influence of the singular points inside the function on the approximation is greatly weakened. Therefore, for functions with smooth boundaries and internal oscillations, the proposed algorithm has a smaller re-solution constant than the original full-data algorithm.
All the results in this paper are based on numerical experiments, and we leave the theoretical proof as a topic for future work. The proposed algorithm has good performance and gives clear conclusions on the settings of various parameters. Therefore, we feel that it is of substantial interest for those who use FEs in practice.
@@ 1.1 Overview of the paper
In Section 2, we outline the algorithm and explain its structure. The specific calculation method of the extension part based on boundary data is given in Section 3. Section 4 systematically tests the key parameters, points out their influencing factors, and gives the optimization setting scheme. In Section 5, we give the numerical test of the final algorithm and compare it with two other fast algorithms.
2 The outline of the boundary interval algorithm
@@
Let the positive integers $m_{\Delta}$ be the number of nodes taken in the boundary interval, and $S_{l},S_{r}$ be the set of corresponding subscripts,
@@
Let $T_{\Delta}$ be a positive interger,
$L_{\Delta}=2\times T_{\Delta}\times m_{\Delta},\quad h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta}.$ (6)
Then let
@@
For $j\in J_{\Delta}$, take
$g(x_{j})=\left\{\begin{aligned} &f(t_{M-m_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t_{-M+j-\frac{L}{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right.$ (8)
Then we calculate $g_{c}(x_{j}),j=1,2,\ldots,L_{\Delta}$ according to the method in section 3. Finally, we will obtain a function $f_{c}$ with a period of $2+\lambda$, where
$\lambda=\frac{L_{\Delta}-2m_{\Delta}+2}{2M},$ (9)
$f_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g_{c}(x_{m_{\Delta}+\ell-M+1}),&\ell>M.\end{aligned}\right.$ (10)
Fig. 2 shows the basic process of the boundary interval algorithm using the function $f(x)=x^{2}$ as an example. Here, in order to make the graph look clearer, we use a relatively small number of nodes. Figure 1 (a) shows the given data (indicated by the black dotted line) and the boundary data selected for the calculation extension (indicated by the red circle). Figure 1 (b) shows the $g_{c}(x_{j})$ calculated from the boundary data. The data represented by the blue circle is the subsequent extension data required. Figure 1 (c) shows the data of $f_{c}$ in one period formed by combining the original data with the calculated extension data. It can be seen from the calculation process that the calculation of the extension part only depends on the boundary interval data.
@@
After obtaining the data in one period, the subsequent algorithm can be implemented through FFT. Therefore, whether the algorithm can achieve satisfactory efficiency depends on the computational complexity of the extension part, that is, the size of the parameters $T_{\Delta}$ and $m_{\Delta}$. We will give their selection schemes through testing in Section 4.
@@ 3 The calculation of $g_{c}$
Including appropriate normalisation, we let
$\phi_{k}(x)=\frac{1}{\sqrt{L_{\Delta}}}e^{\text{i}kx},$ (11)
Assume $g(x_{j})$ is given by (8), then the Fourier extension $g_{c}(x)$ is defined to be
@@
$\gamma=\frac{m_{\Delta}}{n_{\Delta}}.$ (19)
Fig. 3 shows the singular values of $A$ calculated by SVD techniques for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$ in MATLAB. It can be seen that the first approximately $\mathcal{O}(n_{\Delta}/T_{\Delta})$ are close to $1$ and then the singular values rapidly decay until the true singular values are masked by machine precision errors.
(a) $n_{\Delta}=100$, $T_{\Delta}=6$
(b) $m_{\Delta}=100$, $T_{\Delta}=20$
(c) $m_{\Delta}=100$, $n_{\Delta}=100$
The situation of singular values reflects the ill-posedness of (15), regularization needs to be introduced. Referring to the treatment methods in [2], the truncated SVD leas-squares solution to (15) can be given as
@@
where the parameter $\tau$ is usually chosen close to the machine precision.
@@ 4 Numerical investigation and determination of key parameters
The proposed algorithm involves multiple parameters, and their selection has a substantial impact on the performance of the algorithm. This section will test and analyze them and then provide an optimized scheme.
@@ 4.1 Testing the parameter $T_{\Delta}$
Previous research results show that the extension length has a substantial impact on the convergence rate and the resolution power. Generally speaking, a larger extension length can achieve a higher convergence rate but also means a larger resolution constant. For details, refer to [11, 1, 3]. In [1], the authors also point out that the theoretical results may not be witnessed in computations.
For the proposed algorithm, we are more concerned about the influence of $T_{\Delta}$ on the convergence rate, and whether there is any relationship between it and the number of boundary points $m_{\Delta}$, the ratio coefficient $\gamma$, and the frequency of the approximated function. Now we take
$f(t)=\exp(i\pi\omega t)$
to test different parameters. In Fig 4, we give results for the approximation error against $T_{\Delta}$ with different parameters. Here and what follows, the approximation error evaluated as the maximum pointwise error over an equi-sampled grid ten times denser than the one used for construction and the parameter $\tau=1e-14$ (we also tested with $\tau=1e-12,1e-13,1e-15$ and obtained results similar to those shown here.)
@@
Figure 4: The approximation error against $T_{\Delta}$ for various parameters
Three distinct regions are visible in all figures:
  • A region $I_{1}:=\{T_{\Delta}:1<T_{\Delta}<\hat{T}_{1}\}$, where the error decreases rapidly to near the machine accuracy as $T_{\Delta}$ increases.
  • A region $I_{2}:=\{T_{\Delta}:\hat{T}_{1}<T_{\Delta}<\hat{T}_{2}\}$, where the error is maintained close to the machine accuracy.
  • A region $I_{3}:=\{T_{\Delta}>\hat{T}_{2}\}$, where the error rebounds, and the increase of $T_{\Delta}$ will cause the error to increase instead.
It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma$, $M$, and $\omega$. For $\hat{T}_{2}$, we can give an estimate:
$\hat{T}_{2}\sim\frac{M}{\gamma\omega}.$ (21)
For several different values of $\gamma$, we give the approximate values of $\hat{T}_{1}$ (We tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than $135e-$ 13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1.
When the number of nodes $M$ is not large enough, the regions $I_{1}$ and $I_{2}$ are merged into one region. At this time, the error reaches the optimal value near $\hat{T}_{2}$. This phenomenon can be observed from Fig. 5.
@@
What will happen if the $T_{\Delta}$ values selected is less than $\hat{T}_{1}$? Figure 6 shows this. We can see that the error will not decrease after reaching a certain threshold, which cannot be improved by increasing the numbers of $M$ and $m_{\Delta}$. It should be noted that this situation also applies to the Fourier extension algorithm using full data in previous literature. In Figure 7, we show the results of the algorithm using full data. It can be seen that when the value of $T$ is smaller than the corresponding $\hat{T}_{1}$, the method cannot achieve machine accuracy. Similarly, this cannot be improved by increasing the number of $M$.
@@
From the above test results, we can see that the selection of $T_{\Delta}$ is crucial to the approximation accuracy. Only when $T_{\Delta}$ is within the interval $[\hat{T}_{1},\hat{T}_{2}]$, the approximation can obtain satisfactory results. The length and position of this interval are affected by the parameter $\gamma$ and the frequency $\omega$ of the approximated function.

### v4

# Fast Algorithms for Fourier extension based on boundary interval data thanks: The research is partially supported by National Natural Science Foundation of China (No. 12171455, RSF-NSFC 23-41-00002)

@@ Abstract
In this paper, we first propose a new algorithm for the computation of Fourier extension based on data near the two endpoints of the interval, which can obtain an effective Fourier series approximation for non-periodic functions. The algorithm calculates the extended part using the boundary interval data and then combines it with the original data to form the data of the extended function within a period. By testing the key parameters involved, their influence on the algorithm was clarified and an optimization setting scheme for the parameters was proposed. Compared with FFT, the algorithm only needs to increase the computational complexity by a small amount. Then, an improved algorithm for the boundary oscillation function is proposed. By refining the boundary interval grid, the resolution constant of the boundary oscillation function was reduced to approximately $1/4$ of the original method.
@@
Fourier Series, Fourier Extensions, Fourier Continuation, Resolution Constant
@@ 1 Introduction
Fourier series approximation has been widely used in various fields for smooth periodic functions. In many physical, engineering and exploration geophysical problems, such as analyzing periodic signals in electrical engineering, studying periodic motion in mechanics such as the oscillation of a pendulum under certain ideal conditions, or decomposing seismic/electromagnetic signals into different frequency components using Fourier series approximation [17, 19, 20]. The Fourier series approximation offers spectral convergence since more terms are added to the Fourier series, the approximation converges to the actual function in an efficient way in the frequency domain. Moreover, it can be computed numerically via the FFT (Fast Fourier Transform), which is a computationally efficient algorithm. However, the situation changes completely once the periodicity is lost. A jump discontinuity is imposed in the approximation at the domain boundaries, owing to the well-known Gibbs phenomenon. For instance, when we attempt to approximate a function that is not periodic over a given interval using Fourier series, we will observe this Gibbs phenomenon near the boundaries of the interval. Several methods have been developed to ameliorate the ill effects of the Gibbs phenomenon. For example, methods based on Gegenbauer polynomials, which have certain orthogonality properties that can be exploited to obtain better approximate functions [9, 10, 18]. Methods utilizing Padé approximations, which are rational function approximations, can also be used [7, 8]. And Fourier extension (or continuation) methods, which extend the function in a certain way to make it more suitable for Fourier-based approximation, are also among the approaches developed to address this problem [4, 6, 11, 12, 13, 14, 15, 21].
@@
To obtain a solution to the minimization problem (2), a severely ill-conditioned least-squares system has to be solved. In [4], Boyd introduced truncated singular value decomposition (TSVD) to stabilize the calculation process, with an accuracy almost up to machine precision. A convergence analysis by Adcock et al. states that for analytic functions the method is at least superalgebraically convergent [2]. Utilizing randomized algorithms, Lyon provided a fast algorithm with $T=2$ [12]. In [14], Matthysen et al. presented fast algorithms for arbitrary $T$, exploiting it as a specific variant of bandlimited extrapolation. This idea was further developed in [15] to form a fast algorithm for function approximation in 2D domains.
@@
It should be noted that the goal of the extension process is to effectively connect the data at both ends of the function. The following two functions were constructed:
$f_{1}(t)=\left\{\begin{aligned} &-1,&x<-\frac{1}{2},\\ &\tanh\left(\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2},\\ &1,&t>\frac{1}{2},\end{aligned}\right.$ (3)
$f_{2}(t)=\left\{\begin{aligned} &-1,&t<-\frac{1}{2},\\ &\tanh\left(100\tan\pi t\right),&-\frac{1}{2}\leq t\leq\frac{1}{2},\\ &1,&t>\frac{1}{2}.\end{aligned}\right.$
Because the data of the two functions near the endpoints are the same, the same connection data can be used in the intervals $[-T,1)$ and $(1,T]$. An illustration of this process is shown in Fig. 1. This raises the issue of whether the data inside the interval really help in the calculation of the extension. In fact, a very efficient method, known as the FC-Gram method has been developed earlier, which uses only boundary data (projecting it into a basis of orthogonal polynomials) for extension calculations [5]. Recently, this method has also been extended to the approximation of two-dimensional functions[6]. In [16], the authors proposed a method to calculate the Fourier extension using the derivative on the boundary combined with a two-point Taylor interpolation polynomial. The extension part calculated by these two methods are more stable, but neither of them is superalgebraically convergent because polynomials need to be introduced. Judging from the numerical results given in the literature, the number of nodes required by these two methods was significantly higher than that of the previous method when obtaining the same accuracy. In [5], the authors identified a critical limitation in directly applying the original extension algorithm with relatively few nodes in the boundary interval: the inherent difficulty in achieving satisfactory approximation accuracy. Therefore, they turned to precede the FC(SVD) calculation by projecting them into the bases of orthogonal polynomials. However, existing theoretical analyses reveal that the accuracy of the Fourier extension is co-determined by two interdependent parameters –extension length and sampling ratio. In previous studies, researchers tended to increase the sampling ratio to improve accuracy. This is because in the full data algorithm, increasing the extension length does not seem to bring any benefits. In this study, we will point out that by increasing the extension length, the number of nodes required in the boundary interval can be effectively reduced. Specifically, we demonstrate that through optimal parameter combinations, an accurate extension computation can be achieved with a few boundary nodes deployment. The number of nodes required in the boundary interval can be fixed, and is only related to the desired accuracy.
In this paper, we first propose a new fast algorithm for the calculation of the Fourier extension, which has the following traits:
  • By optimizing the parameter settings, only a few nodes in the boundary interval are required to calculate the extension part; Thus, the method can be computed with FFT speed. Unlike the FC-Gram algorithm, the method does not require the introduction of polynomials. When the behavior of the boundary is not significantly better than that of the interior, the proposed algorithm requires significantly fewer nodes than the FC-Gram algorithm because it maintains the superalgebraic convergence property of the original algorithm.
  • Because the calculation of the extension part uses only the data of the small interval near the boundary, the influence of oscillations within the interval on the approximation is significantly weakened. Therefore, for a function that is smooth in the neighborhood of the interval boundary and oscillates within the interval, the proposed algorithm has a smaller resolution constant than that of the original full-data algorithm. This also applies to the FC-Gram algorithm.
Furthermore, for functions that oscillate in the neighborhood of the interval boundary, we propose an improved algorithm based on boundary interval grid refinement, which can significantly reduce the resolution constant of this type of function. The proposed algorithms exhibit good performance and provide clear conclusions regarding the settings of various parameters. Therefore, we believe that it is of substantial interest to those who use FEs in practice.
@@ 1.1 Overview of the paper
In Section 2, we outline the algorithm and explain its structure. The specific calculation method of the extension part based on boundary interval data is presented in Section 3. Section 4 present the computational complexity and resolution constant analysis of the algorithm. In Section 5, we systematically tests the key parameters, identifies their influencing factors, and provide an optimization setting scheme. In Section 6, we present numerical tests of the final algorithm and compare it with three other fast algorithms. Furthermore, for the boundary oscillation function, we propose an improved algorithm based on the boundary interval grid refinement.
2 Outline of the boundary interval algorithm
@@
Let the positive integer $m_{\Delta}$ be the number of nodes taken in the boundary interval, and $S_{l},S_{r}$ be the set of corresponding subscripts,
@@
Let $T_{\Delta}>1$ and
$L_{\Delta}=2\times\left\lceil T_{\Delta}\times(m_{\Delta}-1)\right\rceil,\quad h=\frac{2\pi}{L_{\Delta}},\quad x_{j}=(j-1)h,\quad j=1,2,\ldots,L_{\Delta},$ (6)
[…]
@@
For $j\in J_{\Delta}$, take
$g(x_{j})=\left\{\begin{aligned} &f(t_{M-m_{\Delta}+j}),&j\in J_{\Delta,1},\\ &f(t_{-M+j-{L_{\Delta}}/{2}-1}),&j\in J_{\Delta,2}.\end{aligned}\right.$ (8)
Then we calculate the extension function $g_{c}$ defined on the interval $[0,2\pi]$ of $g$, and obtain its values at $x_{j},j=1,2,\ldots,L_{\Delta}$ according to the method described in Section LABEL:sec31. Finally, we obtain the function $f_{c}$ with a period of $2+\lambda$, where
$\lambda=\frac{\lceil T_{\Delta}-1\rceil\times(m_{\Delta}-1)}{M},$ (9)
$f_{c}(t_{\ell})=\left\{\begin{aligned} &f(t_{\ell}),&|\ell|\leq M,\\ &g_{c}(x_{m_{\Delta}+\ell-M}),&\ell>M.\end{aligned}\right.$ (10)
Fig. 2 shows the basic process of the boundary interval algorithm using the function $f(x)=x^{2}$ as an example. Here, to make the graph clearer, we use a relatively small number of nodes. Fig. 2(a) shows the given data (indicated by the black dotted line) and boundary data selected for the calculation extension (indicated by the red circle). Fig. 2(b) shows the $g_{c}(x_{j})$ calculated from the boundary data. The data represented by blue circles represent the subsequent extension data required. Fig. 2(c) shows the data of $f_{c}$ in one period formed by combining the original data with the calculated extension data. It can be observed from the calculation process that the calculation of the extension part depends only on the boundary interval data.
After obtaining the data in one period, the subsequent algorithm can be implemented using FFT. Therefore, whether the algorithm can achieve satisfactory efficiency depends on the computational complexity of the extension part, that is, the size of the parameters $T_{\Delta}$ and $m_{\Delta}$. We present their selection schemes through testing in Section 5.
@@ 3 The calculation of $g_{c}$
Including the appropriate normalization, we let
$\phi_{k}(x)=\frac{1}{\sqrt{L_{\Delta}}}e^{\text{i}kx},\quad k=-n_{\Delta},\ldots,n_{\Delta},\quad x\in[0,2\pi].$ (11)
Assuming $g(x_{j})$ is given by (8), then the extension function $g_{c}$ of $g$ is defined as
@@
(a) $n_{\Delta}=100$, $T_{\Delta}=6$
(b) $m_{\Delta}=100$, $T_{\Delta}=20$
(c) $m_{\Delta}=100$, $n_{\Delta}=100$
@@
$\gamma_{\Delta}=\frac{m_{\Delta}-1}{n_{\Delta}}.$ (19)
Fig. 3 shows the singular values of $A$ calculated by SVD techniques for various $m_{\Delta}$, $n_{\Delta}$ and $T_{\Delta}$ in MATLAB. It can be observed that the first approximately $\mathcal{O}(n_{\Delta}/T_{\Delta})$ are close to $1$ and then the singular values rapidly decay until the true singular values are masked by machine precision errors.
The situation of singular values reflects the ill-posedness of (15), regularization must be introduced. Referring to the treatment methods in [2], the truncated SVD least-squares solution to (15) can be given as
@@
where parameter $\tau$ is usually chosen to be close to the machine precision.
4 Computational complexity and resolution constant
The computational complexity of the proposed algorithm comprises three parts:
  • Calculation of $c_{\tau}$. The computational effort in this step is mainly concentrated on the singular value decomposition of matrix $A$, which is related to $m_{\Delta}$ and $\gamma$, and can be estimated as
    • If $\gamma_{\Delta}<1$, the computational complexity is
      $\mathcal{O}\left(m_{\Delta}^{2}n_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma_{\Delta}}m_{\Delta}^{3}\right).$ (21)
    • If $\gamma\geq 1$, the computational complexity is
      $\mathcal{O}\left(m_{\Delta}n^{2}_{\Delta}\right)=\mathcal{O}\left(\frac{1}{\gamma_{\Delta}^{2}}m_{\Delta}^{3}\right).$ (22)
    $O\left(L_{\Delta}\log(L_{\Delta})\right).$ (23)
  • Calculation of Fourier coefficients of $f_{c}$. This can be achieved by FFT, and the computational complexity is
    $O\left(M\log(M)\right).$ (24)
Next, we analyze the factors that affect these parameters in detail. First, we provide the following concept for the resolution constant, which comes from the literature [3]:
Definition 4.1.
Let $\left\{F_{M}\right\}_{M\in\mathbb{N}}$ be a sequence of approximations such that $F_{M}(f)$ depends only on the values of $f$ on an equispaced grid of $12M+$ 1 points. Let
$\mathcal{R}(\omega,\delta)=\min\left\{M\in\mathbb{N}:\|\exp(\text{i}\pi\omega t)-F_{M}(\exp(\text{i}\pi\omega t))\|_{\infty}\leq\delta\right\},\quad\omega>0,~{}~{}0<\delta<1,$
then we say that $F_{M}$ has resolution constant $0<r<\infty$ if
$R(\omega,\delta)\sim r\omega,\omega\rightarrow\infty,$ (25)
for any fixed $\delta$.
The resolution constant of the present method is influenced by two aspects:
  • First, we need to consider the approximation ability of $g_{c}$ to the boundary interval data of the function $f$. For the function $f(x)=\exp(\text{i}\pi\omega t)$ on $[-1,1]$, since the calculation of $g_{c}$ is performed on $[0,2\pi]$ after intercepting the boundary interval data, the function corresponding to $g$ in (8) is
    $g(x)=\exp(\text{i}\omega_{\Delta}\pi t),$
    where
    $\omega_{\Delta}=\frac{m_{\Delta}-1}{M}T_{\Delta}\omega.$ (26)
    Obviously, if $g_{c}$ is to be able to effectively approximate $g$, it must satisfy
    $n_{\Delta}\geq\omega_{\Delta}.$ (27)
    Therefore, according to (26), (27) and (19), it can be obtained that the number of nodes $M$ have to satisfy
    $M\geq\gamma_{\Delta}T_{\Delta}\omega.$ (28)
  • Second, it is necessary to consider the number of nodes required for using $f_{c}$ to calculate the overall approximation. In this step we use FFT to approximate a periodic function, according to the basic conclusion of Fourier approximation, this step requires the number of nodes $M$ to satisfy
    $M\geq\omega.$ (29)
Note that in the above analysis process, we consider the case where the internal and boundary frequencies are the same. In the actual calculation process, when facing general functions, it can be seen from the above analysis that if the frequency of the function at the boundary is significantly lower than that of the interior, the number of nodes required for calculation will be greatly reduced. This also applies to the FC-Gram algorithm. We will verify this in the subsequent numerical experiments.
Through (4.1)-(4.3) and the above analysis, we can see that whether the proposed method can have a smaller computational complexity depends on whether there is a smaller $m_{\Delta}$ value that can enable $g_{c}$ to accurately approximate the function in the boundary interval. In the original full-data Fourier extension algorithm, most literatures recommend the use of oversampling mode. The results show that the accuracy of the calculation can be improved with the increase of the sampling ratio
$\gamma=\frac{M}{N}.$
In fact, the increase of $\gamma$ plays a role in reducing the discrete step length on $[-T,T]$, which can also be achieved by increasing $T$, as mentioned in the previous literature on the estimation of the resolution constant of the full-data algorithm [3]:
$r\sim T\gamma.$ (30)
In previous studies on full-data algorithms, researcher tend to increase $\gamma$ because no benefits were observed with increasing $T$. However, for the algorithm proposed in this paper, increasing $T$ seems to be a better choice because it significantly reduces the number of nodes required for the boundary interval. In the next section, we will test these parameters more comprehensively and provide clear results and parameter setting suggestions, so that the algorithm can complete the calculation with lower computational complexity.
@@ 5 Numerical investigation and determination of key parameters
The proposed algorithm involves multiple parameters, and their selection has a substantial influence on the performance of the algorithm. This section tests and analyzes them and then provides an optimized scheme.
@@ 5.1 Testing the parameter $T_{\Delta}$
From the analysis in the previous section, we know that the extension length $T_{\Delta}$ has a substantial influence on the convergence rate and the resolution power. Generally, a larger extension length can achieve a higher convergence rate but also means a larger resolution constant.
In this section, we numerically observe the impact of $T_{\Delta}$ on the performance of the algorithm. Now we take
$f(t)=\exp({\text{i}}\pi\omega t)$
to test the interaction between $T_{\Delta}$ and the other parameters. In Fig 4, we present the results of the approximation error against $T_{\Delta}$ for different parameters. Here and what follows, the approximation error is evaluated as the maximum pointwise error over an equi-sampled grid that is ten times denser than that used for construction and the parameter $\tau=1e-14$ (we also tested with $\tau=1e-12,1e-13,1e-15$ and obtained results similar to those shown here.) Three distinct regions are visible in all graphs in Fig. 4:
  • Region $I_{1}:=\{T_{\Delta}:1<T_{\Delta}<\hat{T}_{1}\}$, where the error decreases rapidly to the machine accuracy as $T_{\Delta}$ increases.
  • Region $I_{2}:=\{T_{\Delta}:\hat{T}_{1}<T_{\Delta}<\hat{T}_{2}\}$, where the error is maintained close to the machine accuracy.
  • Region $I_{3}:=\{T_{\Delta}>\hat{T}_{2}\}$, where the error rebounds, and the increase in $T_{\Delta}$ causes the error to increase instead.
@@
Figure 4: The approximation error against $T_{\Delta}$ for various parameters.
@@
It can be seen that the value of $\hat{T}_{1}$ is influenced by $\gamma_{\Delta}$, but not by $\omega$, $m_{\Delta}$, and $M$, while $\hat{T}_{2}$ is influenced by $\gamma_{\Delta}$, $M$, and $\omega$. For several different values of $\gamma_{\Delta}$, we give the approximate values of $\hat{T}_{1}$ (we tested the functions of $\omega$ from $1$ to $50$, and took the mean of the corresponding $T_{\Delta}$ when their error was less than $131e-$ 13 for the first time as the estimate of $\hat{T}_{1}$), which are listed in Table 1. For $\hat{T}_{2}$, we can estimate:
$\hat{T}_{2}\sim\frac{M}{\gamma_{\Delta}\omega}.$ (31)
This is consistent with the result obtained in equation (28).
@@
When the number of nodes $M$ is not sufficiently large, regions $I_{1}$ and $I_{2}$ are merged into a single region. At this time, the error reached the optimal value near $\hat{T}_{2}$. This phenomenon is shown in Fig. 5.
What will happen if the $T_{\Delta}$ value selected is less than $\hat{T}_{1}$? Figure 6 shows this. We can see that the error will not decrease after reaching a certain threshold, which cannot be improved by increasing the numbers of $M$ and $m_{\Delta}$. It should be noted that this situation also applies to the Fourier extension algorithm using the full data in previous literatures. In Fig. 7, we show the results of the algorithm using full data. It can be seen that when $\gamma$ is fixed and the value of $T$ is smaller than the corresponding $\hat{T}_{1}$, the method cannot achieve machine accuracy. Similarly, this cannot be improved by increasing the number of $M$.
@@
From the test results above, we can see that when $\gamma_{\Delta}$ is fixed, the selection of $T_{\Delta}$ is crucial to the approximation accuracy. The approximation can obtain satisfactory results only when $T_{\Delta}$ is within the interval $[\hat{T}_{1},\hat{T}_{2}]$. The left endpoint $\hat{T}_{1}$ of the interval is determined by parameter $\gamma_{\Delta}$, whereas the right endpoint $\hat{T}_{2}$ is affected by parameters $\gamma_{\Delta}$, $\omega$ and $M$.